OpenStudy (anonymous):

for all integers r and s, let r#s=4r-3s. which of the following must be true? I. 4#3=7 II. (1#0)#2=1#(0#2) III. r#s=s#r

hartnn (hartnn):

r#s=4r-3s 4#s means r=4, s=3 = 4*4 - 3*3 is it = 7 ?

OpenStudy (anonymous):

do you mean the answer is only I ?? @hartnn

hartnn (hartnn):

1 is true i didn't check II. did u?

hartnn (hartnn):

III. is not true

hartnn (hartnn):

1#0 =4 4#2 = 10 0#2 =-6 1#-6 = 4+18 = 22 so II is not true

hartnn (hartnn):

only I is true...unless i did some careless calculation mistake :P verify it...

OpenStudy (anonymous):

can i know whats the function of # ?

hartnn (hartnn):

its already defined! \( r\#s=4r-3s \)

hartnn (hartnn):

in words, "multiply first number by 4, 2nd number by 3 and take the difference"

OpenStudy (anonymous):

ohh i got it now thank you @hartnn

hartnn (hartnn):

welcome ^_^